OpenStudy (anonymous):

Henry rolls 2 number cubes numbered 1 through 6 while playing his favorite board game. He will get a second turn if he rolls a sum that is an even number less than 10. What are Henry's chances of getting a second turn when he rolls the number cubes?

OpenStudy (anonymous):

Here are the options: A. \[\frac{ 7 }{ 18 }\] B. \[\frac{ 11 }{ 18 }\] C.\[\frac{ 5 }{ 36 }\] D.\[\frac{ 17 }{ 36 }\]

OpenStudy (yanasidlinskiy):

Start by calculating how many ways Henry can get an 'even sum'